OpenStudy (fellowroot):

My question is about part B. Is it slowing down or speeding up? Some of my classmates said that if both your velocity and acceleration are negative then its actually speeding up. I disagree and say that the only function you need to look at is the acceleration one. If at t=4 the acceleration is negative then it should be slowing down. Please help. thanks.

OpenStudy (fellowroot):

pic

OpenStudy (anonymous):

Hi Fellowroot...since your question is constraining you to a number line, as long as you dictate that your forward direction on the number line is positive, then if anywhere on that number line you have a negative acceleration, you are slowing down... Think about it this way... At the very beginning of your analysis of your movement along this number line, at time equal to ZERO, you would have a positive forward velocity of 12m/s... at t=1, your velocity is +6 m/s at t=2, your velocity is zero at t=3 your velocity is -15 m/s and as you stated at t=4, your velocity is -36 m/s but with your reference frame being the number line itself for this whole analysis... So between time zero and time=4, you slowed down(negative acceleration), came to a brief stop, and started moving faster per second in the opposite direction... If you analyze your problem with these factors in mind, unfortunately neither you or your fellow students have the entire picture correct. In actuality, you would be accelerating, but always keeping your reference frame in mind, you would be accelerating in the "opposite" direction. Hope this helps! :o)

OpenStudy (anonymous):

I guess what I am saying is that taking a snippet of what is really going on only at time=4 can get confusing...look at the whole picture and things become much clearer!

OpenStudy (fellowroot):

Hey remember me from like a few years ago. how is school? what degree are you going for?

OpenStudy (anonymous):

Hi fellowroot... a physics degree:o) did my answer help at all?